OpenStudy (anonymous):

What is the value of –3 • (7 + (–3)2) ÷ 3?

OpenStudy (anonymous):

-1

OpenStudy (shaik0124):

-3.(7+(-6))/3 =-3.(7-6)/3

OpenStudy (anonymous):

whats the answer ?

OpenStudy (shaik0124):

-1

OpenStudy (anonymous):

What is the solution to 8 • g = –32?

OpenStudy (shaik0124):

g=-32/8

OpenStudy (anonymous):

what is the answer your not giving me it

OpenStudy (shaik0124):

-4

OpenStudy (anonymous):

Which expression can be written as 4 • (3 + 8)?

OpenStudy (shaik0124):

do u know tables and divisions

OpenStudy (anonymous):

no im stupid thats y im on this thing

OpenStudy (shaik0124):

ok u will b better one day but don't give up or hate that subject rather love it it will give u more n more

OpenStudy (anonymous):

ok can u give me the answer now

OpenStudy (shaik0124):

which answer

OpenStudy (shaik0124):

44